OpenStudy (anonymous):

find the vertex: 4x^2-y+6=8x

sam (.sam.):

Do you know what is vertex?

sam (.sam.):

A vertex can be said as corners of intersections or maximum where the line switches directions.

sam (.sam.):

So to find the maximum point you have to find the maximum or minimum point. You can do this by completing the square.

sam (.sam.):

4x^2-y+6=8x Arrange the equation to quadratic equation form, \[y=4 x^2-8 x+6\]

sam (.sam.):

4(x^2-2x)+6 4(x-1)^2-1(4)+6 4(x-1)^2+2

sam (.sam.):

4(x-1)^2+2 You can determine the coordinates from here, (x-1)=0 x=1 ------------------------------------------- y=2 So, (1,2)

OpenStudy (anonymous):

formula for vertic x=-b/2a then put x in equation to find y x,y will be vertix

OpenStudy (anonymous):

first arrange equ like y=ax^2+bx+constant

OpenStudy (anonymous):

can you solve it now?

OpenStudy (anonymous):

tell me so that i solve it for you if u dont

OpenStudy (anonymous):

i solved it